6. [-/1 Points] DETAILS LARSON8 9.2.036. Find the sum of the convergent series. 12 n(n +2) n=1
Answer #1

12 L_pn12) m4 A 12 3 12 su 1) + (कर 122 A (n-t2) t Br put nzo 12 - A.2 A26 put n=-2 122-28 32 -6 6 6 > mi 2 n . 12 (12) 12 -26 K+2 फिट) ae bw Ton- E+ -(ra) = 1++ -ला- - (1-2) + (14) + (3-5-4-1) n-3 - n=2 m ल n+2 [++J= ME intana य 11 =bm C 1-2-3 6 (-+